OpenStudy (anonymous):

can anyone explain how the equation got 13/4?

OpenStudy (anonymous):

The displacement (in meter) of an object moving in a straigt line is given by s=1+2t+1/4(t^2) a) find the avarage veolocity over each time period

OpenStudy (anonymous):

the answer says \[f(x)=\frac{ s(1+h)-s(1) }{ (1+h)-1 }=\frac{ 1+2(1+h)+(1+h)^2/4-13/4 }{ h }\]

OpenStudy (anonymous):

@Twis7ed

OpenStudy (anonymous):

why they got -13/4 in the end?

OpenStudy (anonymous):

That is the value of s(1)

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

Do you understand the first step they took? The s(1+h)-s(1)/(1+h)-1?

OpenStudy (anonymous):

Yes, I understand that part.

OpenStudy (anonymous):

Ok, then the 13/4 is the value of the s(1) part of the equation

OpenStudy (anonymous):

Since you get 13/4 when you plug in 1 for the x in your s(x) function

OpenStudy (anonymous):

which equation?

OpenStudy (anonymous):

s(1+h)−s(1)/(1+h)−1

OpenStudy (anonymous):

oh I see.... but i am still confused about -13/4

OpenStudy (anonymous):

In the equation s(1+h)−s(1)/(1+h)−1, the part that says s(1) is basically the value of the equation s(x) which they gave you, when x=1 so you can just plug in 1 for x and solve and you end up getting 13/4

OpenStudy (anonymous):

ohh I see. so the equation told me is equation for any velocity problem?

OpenStudy (anonymous):

the equation ( you told me)